OpenStudy (anonymous):

What is the sum of the first 13 terms in the Fibonacci Sequence? Term 1 2 3 4 5 6 7 8 9 Fibonacci Number 1 1 2 3 5 8 13 21 34

OpenStudy (mertsj):

Could you write the first 13 terms and then add them up?

OpenStudy (asnaseer):

one way of doing this is to just work out the 1st 13 terms and then just add them up.

OpenStudy (mertsj):

You already have 9 of them in your post.

OpenStudy (asnaseer):

you have been given the 1st 9 terms above.

myininaya (myininaya):

Thats a confusing question I think. Some people start this sequence with 0 and some start it with 1.

OpenStudy (asnaseer):

@Mertsj - we seem to be experiencing ESP here ;-)

OpenStudy (anonymous):

i get 609

OpenStudy (anonymous):

12 is 144 and 13 is 233

OpenStudy (anonymous):

but i cheated http://www.wolframalpha.com/input/?i=sum+n+%3D+1to+13+fibonnaci

OpenStudy (anonymous):

thank u satellite

OpenStudy (anonymous):

how did u fined that site

myininaya (myininaya):

that site is popular

OpenStudy (mr.math):

Lets find the sum of the first 100 Fibonacci numbers.

OpenStudy (asnaseer):

@sunsetlove - I would still advise you to try and work this out yourself so that you understand it better. do you know how the Fib series is constructed?

OpenStudy (anonymous):

no i dont

myininaya (myininaya):

1 1 1+1=2 1+2=3 2+3=5 3+5=8 5+8=13 8+13=21 13+21=34

OpenStudy (asnaseer):

so you have the first two terms as 1, 1 the next one is the sum of these two terms, so 1+1=2 now we have 1, 1, 2 next term is sum of last two terms, so 1+2=3 now we have 1, 1, 2, 3 next term is sum of last two terms, so 2+3=5 now we have 1, 1, 2, 3, 5 etc

OpenStudy (anonymous):

i tryed that it confused me so 21+36 is next right ?

OpenStudy (asnaseer):

21+34

OpenStudy (asnaseer):

the last two terms in the sequence given to you are 21 and 34, so next term will be the sum of these last two terms: 21+34=55

OpenStudy (asnaseer):

so 10th term is 55 which means 11th term is 55+34=89

OpenStudy (asnaseer):

do you see the pattern now?

OpenStudy (anonymous):

and the 11th is 34+89=123 right

OpenStudy (asnaseer):

remember, to get the next term, you add the previous two terms so we know: 10th term is 55 11th term is 89 so 12th term = 55+89=144

OpenStudy (anonymous):

owww ok i get it

OpenStudy (asnaseer):

good - you are in a much stronger position now - mathematically speaking ;-)

myininaya (myininaya):

\[f_n=f_{n-1}+f_{n-2} , n >2 ,f_1 =1 ,f_2=1\]

OpenStudy (asnaseer):

this pattern of numbers is actually quite interesting as it arises a lot in nature - you may wish to investigate if you are interested...

OpenStudy (anonymous):

so 13 would be 89+=233

OpenStudy (asnaseer):

yup

OpenStudy (anonymous):

im anser would be 233 ?

OpenStudy (asnaseer):

no - you need to add up all the values in the series.

OpenStudy (asnaseer):

up to the 13th one

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

so sum=1+1+2+3+5+8+...+89+233

OpenStudy (mertsj):

1+1+2+3+5+8+13+21+34+55+89+144+233=609 and that is the sum of the first 13 terms of the sequence

OpenStudy (mertsj):